OpenStudy (zeesbrat3):

A particle moves on a line away from its initial position so that after t hours it is s = 2t2 +3t miles from its initial position. Find the average velocity of the particle over the interval [1, 4].

OpenStudy (zeesbrat3):

@Hero @nincompoop @abb0t @Whitemonsterbunny17 @freckles @ganeshie8 @Miracrown @vera_ewing @jim_thompson5910

OpenStudy (zeesbrat3):

any ideas?

OpenStudy (zeesbrat3):

help please..

ganeshie8 (ganeshie8):

average rate of change (average velocity) between two points is same as the "slope" of the secant line connecting those two points

OpenStudy (zeesbrat3):

So find the derivative and set it equal to 0?

ganeshie8 (ganeshie8):

You're given \[s(t) = 2t^2 +3t\] For average velocity in the interval \([1,4]\), you simply find the slope between points \((1,s(1))\) and \((4,s(4))\) : \[\dfrac{s(4)-s(1)}{4-1}\]

OpenStudy (zeesbrat3):

mean value theorem..

ganeshie8 (ganeshie8):

Easy, thats just the slope formula!

OpenStudy (zeesbrat3):

\[s(4) = 2(4)^2 + 3(4) = 32 + 12 = 44\] \[s(1) = 2 + 3 = 5\] \[\frac{ 44 - 5 }{ 4 -1 } = \frac{ 39 }{ 3 } = 13\]

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (zeesbrat3):

That's it?

ganeshie8 (ganeshie8):

Yep.

OpenStudy (zeesbrat3):

Wow! Thank you, as always :)

ganeshie8 (ganeshie8):

np

OpenStudy (zeesbrat3):

@ganeshie8 what would be the unit? miles/hour^2?

ganeshie8 (ganeshie8):

how do you measure velocity ?

OpenStudy (zeesbrat3):

\[\frac{ m }{ s^2 }\]

ganeshie8 (ganeshie8):

Really ?

OpenStudy (zeesbrat3):

I was actually good at physics lol

OpenStudy (zeesbrat3):

I would normally say that but this problem has hours and miles

OpenStudy (zeesbrat3):

i lied its just \[\frac{ m }{ s }\]

OpenStudy (zeesbrat3):

\[\frac{ m }{ s^2 }\] is acceleration

ganeshie8 (ganeshie8):

\[\text{average rate of change}=\dfrac{s(4)-s(1)}{4-1}\] Notice that the units for top is \(miles\) and the units for bottom is \(hours\). so the unit for average rate of change is \(miles/hour\)

OpenStudy (zeesbrat3):

So, just switch from metric to american

ganeshie8 (ganeshie8):

average velocity of the particle over the interval [1, 4] is 13 miles/hour

OpenStudy (zeesbrat3):

Okay, thank you :)